OpenStudy (anonymous):

Can someone help me complete the integral of: sec(t) - cos(t)? Is it in an integral table? or is it done through some sort of substitution?

OpenStudy (anonymous):

integral table

OpenStudy (anonymous):

okay. I'll check it out then! Thanks :)

OpenStudy (anonymous):

log(sec t+tant)-sint+c

OpenStudy (anonymous):

One other thing: \[\int sec(t) - cos(t)\ dt = \int sec(t) \ dt- \int cos(t)\ dt\]

OpenStudy (anonymous):

So what you are really asking is what is the integral of sec since the other integral should be easy for you.

OpenStudy (anonymous):

Alchemista, thank you for the framing this from a cleaner perspective. I should have known that it was just int of sec!

OpenStudy (anonymous):

And now I realize that if I didn't have an int table, I should have multiplied the sec by (sect + tant)/(sect + tant). and then integrate.

OpenStudy (anonymous):

though i think one will still need a table....